OpenStudy (lgbasallote):

Prove that the sum of x squares is \[\frac{3x^3 + 2x^2 + x}6\]

OpenStudy (lgbasallote):

is it possible to prove this using some method other than induction?

OpenStudy (anonymous):

yes there is

OpenStudy (lgbasallote):

what would it be?

OpenStudy (anonymous):

Ha ha ha ha.. Hey lgba how are you???

OpenStudy (anonymous):

How you posted two questions lgba ??

Parth (parthkohli):

He has magical powers. lol

OpenStudy (anonymous):

Yeah I know..

OpenStudy (anonymous):

Hey parth how are you??

Parth (parthkohli):

Hi gsRttN, I'm great and you?

OpenStudy (anonymous):

Yeah I am also fine.. How is your summer school going ??

Parth (parthkohli):

Summer school?

OpenStudy (anonymous):

Yeah your school name...

OpenStudy (anonymous):

The solution is very large..

OpenStudy (anonymous):

It should be \[\frac{2x^3 + 3x^2 + x}6\] not\[\frac{3x^3 + 2x^2 + x}6\]@ihbasallote

Parth (parthkohli):

How do you know, gsRttN?

Parth (parthkohli):

@gsRttN

OpenStudy (anonymous):

Can't I know this??? @ParthKohli

OpenStudy (anonymous):

The solution is very large to be written here, so I have given the link...

Parth (parthkohli):

lol hi waterineyes.

Parth (parthkohli):

@gsRttN: why did you make a new account?

OpenStudy (anonymous):

If you don't like induction (I don't either, sometimes there is no choice) you can make use of the following: Sum to n [(i + 1)k - i^k] = (n+1)^k -1 Setting k to 2, 3 etc will produce formulas (with some algebra, of course) for sums of first, second, third powers....

OpenStudy (lgbasallote):

...isn't that induction too?

OpenStudy (anonymous):

U know I'm fond of explicit constructions, right?

OpenStudy (anonymous):

So to do with k = 2 -> (i+1)^2 = 2^2 + 3^2 +.... +(n-1)^2 + n^2 + (n+1)^2 i^2=1^2+ 2^2 + ........ + n^2 and subtract -> -1 + (n+1)^2 which is twice the usual formula for sum of integers 1 to n.

OpenStudy (anonymous):

However, induction is easier, just less informative....

OpenStudy (anonymous):

you can prove one from the previous one

OpenStudy (anonymous):

that is to say, if you know \[\sum_{k=1}^nk\] then you know \[\sum_{k=1}^nk^2\] by algebra and also if you know \[\sum_{k=1}^nk^2\] then algebra gives you \[\sum_{k=1}^nk^3\] etc would be happy to show you when you are here

OpenStudy (anonymous):

btw if you know the answer is a polynomial you don't even need induction, just compute for three values and you are done

OpenStudy (lgbasallote):

i'm confused...isn't proving using summation considered induction?

OpenStudy (anonymous):

not at all, and it usually is totally unnecessary inductions proofs for summation formulas are trivial, since reducing to the previous case is nothing more than saying "sum up to the second to last term"

OpenStudy (lgbasallote):

is it possible to prove this without using those summations at all? honestly, i'm very bad with summation

OpenStudy (anonymous):

but in any case let me outline how you can find \[\sum_{k=1}^nk^2\] if you know \[\sum_{k=1}^nk=\frac{n(n+1)}{n}\]

OpenStudy (anonymous):

Sure, you can do it like I said but it is a lot of writing........

OpenStudy (anonymous):

oh not at all

OpenStudy (anonymous):

With summation symbols, it is much shorter

OpenStudy (anonymous):

But usually gives younger students a headache:-)

OpenStudy (lgbasallote):

hmm...maybe i should try learning summation first....

OpenStudy (anonymous):

ok i have a recommendation for you then work through this

OpenStudy (anonymous):

the idea is to work from the previous case, so for example if you know \[\sum_{k=1}^nk=\frac{n(n+1)}{2}\] then you consider \[n^3=\sum_{k=1}^nk^3-(k-1)^3\] do some algebra, and see that you get an expression in terms only of \(k^2\) and \(k\) then do a but more algebra to solve for the sum you want

OpenStudy (anonymous):

and yes, you should learn summation first